OpenStudy (anonymous):

Use properties of logarithms to condense the logarithmic expression. Write the expression as a single logarithm whose coefficient is 1. Where possible, evaluate logarithmic expressions without using a calculator. log(x)+log(x^2-9)-log15-log(x+3)

zepdrix (zepdrix):

Hmm this one shouldn't be too bad. It's pretty much what we did in the last problem, just in reverse.

OpenStudy (anonymous):

yay :)

zepdrix (zepdrix):

\[\Large\rm \color{orangered}{\log(x)+\log(x^2-9)}-\log15-\log(x+3)\]We'll condense these first two terms down using a rule of logs:\[\Large\rm \log(a)+\log(b)=\log(ab)\]

OpenStudy (anonymous):

log(x(x^2-9))

zepdrix (zepdrix):

When you bring the x into the log, make sure it multiplies the `entire` x^2-9, not just the x^2. So adding another set of brackets might help.\[\Large\rm \color{orangered}{\log\left[x(x^2-9)\right]}-\log15-\log(x+3)\]Oh you're too quick for me c: very nice!

zepdrix (zepdrix):

Then we'll apply our other rule of logs,\[\Large\rm \log(a)-\log(b)=\log\left(\frac{a}{b}\right)\] to these two terms:\[\Large\rm \color{#CC0033}{\log\left[x(x^2-9)\right]-\log15}-\log(x+3)\]

OpenStudy (anonymous):

help :(( hahahahaha

OpenStudy (anonymous):

I'm sorry

zepdrix (zepdrix):

you -_-

zepdrix (zepdrix):

So the rule is telling us we can condense the two logs down to a single log, and divide by the 15, yes?\[\Large\rm \color{#CC0033}{\log\left[\frac{x(x^2-9)}{15}\right]}-\log(x+3)\]

OpenStudy (anonymous):

is it log(x(X^2-9)-log15/log(x+3)

zepdrix (zepdrix):

hmm no :o try to follow along missy!

OpenStudy (anonymous):

sorry that sent late!

OpenStudy (anonymous):

but okay im following you again continue :)

zepdrix (zepdrix):

Oh laggy? :o that's no fun.

zepdrix (zepdrix):

So we'll apply that same division rule to the last two terms.

zepdrix (zepdrix):

We'll combine them into one log, throwing the second term into the denominator.

zepdrix (zepdrix):

\[\Large\rm \log\left[\frac{x(x^2-9)}{15(x+3)}\right]\]Something like that, yah? :d

OpenStudy (anonymous):

yes yes !

OpenStudy (anonymous):

is that the final answer?

zepdrix (zepdrix):

So we've successfully smashed it down into a single log using our fancy rules. And it doesn't have any coefficient showing in front, which means the coefficient is 1. So yay! Looks like we did everything they asked for.

OpenStudy (anonymous):

can you type that answer into this engine? I don't think I'm entering it how I'm suppose to

zepdrix (zepdrix):

\[\Large\rm \log\left[\frac{x(x^2-9)}{15(x+3)}\right]\] =log((x(x^2-9))/(15(x+3)))

OpenStudy (anonymous):

its asking if anything inside the log can be factored or cancelled

zepdrix (zepdrix):

oh oh oh. my bad.

zepdrix (zepdrix):

We have the `difference of squares` in the numerator.\[\Large\rm x^2-9=x^2-3^2\]Difference of squares can be factored into the product of conjugates:\[\Large\rm a^2-b^2=(a-b)(a+b)\]

zepdrix (zepdrix):

\[\Large\rm x^2-3^2=(x-3)(x+3)\]

zepdrix (zepdrix):

And then it looks like we can cancel out our (x+3)'s, yes?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\Large\rm \log\left[\frac{x(x-3)\cancel{(x+3)}}{15\cancel{(x+3)}}\right]\]

zepdrix (zepdrix):

= log((x(x-3))/15) Hopefully that works :o

OpenStudy (anonymous):

so is the final answer than log(x(x-3)/15) ?

zepdrix (zepdrix):

Yah that looks better. I guess I put more brackets than was needed.

OpenStudy (anonymous):

well thank you so so much :)

OpenStudy (anonymous):

youre the best!